OpenStudy (anonymous):

prove the conjunction identity sin(pi/2-x)=sin (pi/2+x)

OpenStudy (anonymous):

guess you are supposed to use the addition angle formula on the right and the subtraction angle formula on the left and see that you get the same thing

OpenStudy (anonymous):

which ought to be \(\cos(x)\)

OpenStudy (anonymous):

not sure that helped, if you are still confused let me know

OpenStudy (anonymous):

well my teacher told us not to do it that way so I do not know what other way I can prove that the two are equal

OpenStudy (anonymous):

not to use the fact they they were both cosine?

OpenStudy (anonymous):

yes. He wanted us to prove that they were equal by changing on side of the equation to match the other side

OpenStudy (anonymous):

i can show you why they are obviously the same

OpenStudy (anonymous):



OpenStudy (anonymous):

another thing you could try (requires knowing that \(\sin(x)=-\sin(x+\pi)\) is to right the left hand side as \[\sin(\frac{\pi}{2}-x)=-\sin(x-\frac{\pi}{2})\]

OpenStudy (anonymous):

that because sine is odd

OpenStudy (anonymous):

lol *write the left hand side ..."

OpenStudy (anonymous):

then add \(\pi\) to the input, change the sign again and get \(\sin(x+\frac{\pi}{2})\)

OpenStudy (anonymous):

ok that helps thank you